OpenStudy (anonymous):

for which equation is x=-1/2 a solution? a. 7-4x=-9 b. 9x+2=7 c. 6+8x=-10 d. 10x-3=-8

OpenStudy (anonymous):

Plug in x for each equation until the solution is found 7-4(-1/2)=-9 7+2=-9 \[9\neq-9\] 9(-1/2)+2=7 -4.5+2=7 \[-2.5\neq7\] 6+8(-1/2)=-10 6-4=-10 \[2\neq-10\] 10(-1/2)-3=-8 -5-3=-8 \[-8=-8\] x=-1/2 is a solution for equation d.

OpenStudy (anonymous):

thanks for the huge helped! you rock :)